# Belief Bias Identification

### v2

@@ Abstract
This paper proposes a unified theoretical model to identify and test a comprehensive set of probabilistic updating biases within a single framework. The model achieves separate identification by focusing on the updating of belief distributions, rather than classic point-belief measurements. Testing the model in a laboratory experiment reveals significant heterogeneity at the individual level: All tested biases are present, and each participant exhibits at least one identifiable bias. Notably, motivated-belief biases (optimism and pessimism) and sequence-related biases (gambler’s fallacy and hot hand fallacy) are identified as key drivers of biased inference. Moreover, at the population level, base rate neglect emerges as a persistent influence. This study contributes to the belief-updating literature by providing a methodological toolkit for researchers examining links between different conflicting biases, or exploring connections between updating biases and other behavioral phenomena.
00footnotetext: For their invaluable support, I would like to thank Elias Tsakas, Thomas Meissner and Matthias Wibral. For helpful comments, I thank Mohammed Abdellaoui, En Hua Hu, Peiran Jiao, Emmanuel Kemel, Diego Marino-Fages, Arno Riedl, David Rojo Arjona, Michael Thaler and seminar participants at IMEBESS in Riga, ASFEE in Grenoble, Behavioural Game Theory Workshop in Norwich, Cognitive Foundations of Decision Making in Rabat, and Maastricht University. I acknowledge funding from the Maastricht Graduate School of Business and Economics (GSBE) for primary data collection 2023 No. G.23.5168.
@@ 1 Introduction
Over the last decades, there have been substantial efforts to identify and categorize belief-updating biases, both within Economics and Psychology (Edwards,, 1968; Tversky and Kahneman,, 1974; Grether,, 1980). Scholars have typically studied these biases separately. This has made it more challenging to tell if someone’s updating behavior can be attributed to one bias or another. Despite recent shifts towards a more comprehensive structure of these biases (Benjamin,, 2019; Bordalo et al., 2023a, ; Bordalo et al., 2023b, ; Stango and Zinman,, 2023), there is still a lack of a unifying testable model that can tell multiple conflicting biases apart.
The essential reason why such a model would be desirable is that belief-updating biases could be observationally equivalent. For example, imagine a person who consistently updates her beliefs “too much” in the face of new information. Traditional models are able to tell whether this behavior is due to this person neglecting her prior beliefs (base rate neglect), or interpreting too much out of the information she receives (overinference). However, other potential biases could also explain her updating behavior. If she has preferences over different outcomes, is she updating her beliefs too much because she generally overinterprets information, or because she cares about the outcome (optimism/pessimism)? Or may she instead be jumping to conclusions because the information she receives is confirming her previous beliefs (confirmation bias)? Could it also be that she has very little doubt about the outcome that she believes to be more likely to happen (overconfidence)? By excluding some of these biases from the model, we may wrongly attribute her behavior to a different bias than the one(s) she is actually exhibiting. This implies that some of the biases which happen to be present in the literature may only seem to be relevant because we lack a more complete model. For instance, base-rate neglect might appear to be driving behavior because, say, confirmation bias and overconfidence are unaccounted for.
@@
Addressing confounded belief-updating biases is important for two related purposes. First, identifying these biases separately is relevant because the actions that people may take under different updating biases could be completely different even if they lead to the same posterior. For example, someone suffering from confirmation bias may keep voting for the same political party after receiving social media information that aligns with her previous beliefs, while someone who simply neglects prior information might be more likely to change her mind if reminded of the actual base rate. Secondly, once confoundedness is taken care of, such a model would be able to shed light onto which specific biases are more common and which biases drive inference the most.
Integrating multiple belief biases into one testable model poses one important methodological challenge. Namely, that the way we usually measure beliefs may not suffice to separably identify a full range of potentially confounded biases. The vast majority of the literature measures point beliefs to study biased updating (Benjamin,, 2019). This approach, while simple, is not sufficient to distinguish multiple observationally-equivalent biases. Because point beliefs offer reduced flexibility to tell one updating pattern apart from another, I propose an alternative framework to address these challenges.
@@
This paper constructs a theoretical model that allows to separably identify multiple conflicting belief-updating biases, using belief distributions [^1] While Jiao et al., (2020) utilize the idea of using belief distributions to simultaneously incorporate belief biases, their model is unsuitable for probabilistic biases (relies on normally distributed beliefs), incorporates a reduced range of biases and is not tested in an experimental setting. to measure prior and posterior beliefs. One can understand belief distributions as an individual holding beliefs over multiple outcomes, or being uncertain about her own beliefs when there are only two states (in a similar sense to cognitive uncertainty (Enke and Graeber,, 2023)). Either way, the central idea is that belief distributions provide a more precise definition of belief biases, offering greater leeway to distinguish between different updating patterns.
@@
The theoretical framework I propose adopts an approach where an agent observes a series of Bernoulli trials. The agent’s prior beliefs are assumed to follow a beta distribution and are updated via Bayesian inference. Since the beta distribution is conjugate to the binomial distribution, the posterior of the Bayesian agent will also result in a beta distribution [^2] This updating setting is comparable to the one employed in the ambiguity aversion literature (see Abdellaoui et al., (2021) as an example) or cognitive uncertainty (Enke and Graeber,, 2023).. To integrate different biases, the model accounts for deviations from Bayesian updating, utilizing distorted likelihoods and priors. These distortions yield non-Bayesian posterior beta distributions, allowing for the identification of core biases such as over/under inference (Khaw et al.,, 2021; Ba et al.,, 2022; Augenblick et al.,, 2024) or base rate neglect/over use (Benjamin et al.,, 2019; Bucher and Glimcher,, 2022; Enke et al., 2023a, ). It also explores asymmetries in reacting to good and bad news in the observed data generating process, alongside preference-based biases (Eil and Rao,, 2011; Zimmermann,, 2020; Möbius et al.,, 2022). Further, it incorporates confirmation bias (Rabin and Schrag,, 1999; Charness and Dave,, 2017; Zhenxun,, 2024), sequence-related biases like the hot hand fallacy and gambler’s fallacy (Rabin,, 2002; Rabin and Vayanos,, 2010; Prat-Carrabin et al.,, 2024), and confidence biases (Serra-Garcia and Gneezy,, 2021; Huffman et al.,, 2022; Enke et al., 2023b, ) that affect the variance of the agent’s posterior distribution, all within the same framework.
@@
The model is tested with help of a novel laboratory experiment. Participants solved a series of tasks where they had to guess the percentage of red balls in a selected urn from a pool of 99 urns, each containing varying proportions of red and blue balls. Participants received information signals via ball draws from the selected urn, forming and reporting their belief beta distributions twice. Then they reported their updated beliefs after observing another sequence of independently drawn signals from the same urn.
The results of the study support the idea that the inclusion of a more complete set of biases mitigates the artificial presence of some biases in more reduced models. At the population level, both overinference and base rate neglect significantly affect updating behavior, when these are the only biases that the model accounts for. However, when the model encompasses a broader spectrum of biases, the impact of overinference vanishes, leaving base rate neglect as the sole remaining persistent bias.
Nonetheless, if people exhibit different biases, findings at the aggregate level might not be able to capture the importance of each bias, as many of them will be averaged out across individuals. In other words, due to the presence of individual heterogeneity, [^3] The idea of individual heterogeneity being a key determinant of biased belief updating has been recently discussed by Khaw et al., (2021) and Alós-Ferrer and Garagnani, (2023); supporting the idea that while average reports might look almost Bayesian or noisy at the population level, an individual analysis reveals that there are different but systematic deviations among individuals. the experiment arguably offers even more insightful findings when applied at the individual level. First, when only a reduced amount of biases is accounted for, the model is unable to identify biased inference in some subjects who would otherwise exhibit a clear bias. In fact, when the full array of biases is factored in, the model shows that all individuals exhibit at least one specific bias; and more importantly, all tested biases are present to a certain extent. Finally, despite the presence of individual heterogeneity, motivated-belief biases (optimism and pessimism) and sequence-related biases (gambler’s fallacy and hot hand fallacy) appear to be the main drivers of biased inference. On the other end, the presence of confirmatory biases (confirmation and disconfirmation bias) is scarce.
This paper does not only contribute to the literature of belief updating as a compass to gauge the intensity of various belief biases, but also aims to be a toolkit for researchers who identify belief-updating biases in settings where these could be potentially confounded. In this direction, there are three ways in which this paper contributes to the existing literature.
@@
First, there has been quite some recent work trying to trace links between different updating biases. For example, Heger and Papageorge, (2018) and Gneezy et al., (2023) study how wishful thinking (optimism) can affect overconfidence; Charness and Dave, (2017) and Zhenxun, (2024, 2023) try to separate behavior stemming from motivated beliefs and unmotivated confirmation bias; and Aydogan et al., (2022) build on Rabin and Schrag, (1999) to distinguish conservatism from confirmation bias. In the absence of an overarching model that encompasses a wide array of biases, findings can be confounded and results about such links could therefore be biased. For example, bias A may appear to be linked to bias B, only because a third bias C is missing.
@@
Second, many researchers search for belief-based explanations of behavioral phenomena, when conflicting biases come into play. For example, political polarization has been separately explained from the standpoint of overconfidence (Ortoleva and Snowberg,, 2015), and confirmation bias (Del Vicario et al.,, 2017). In the financial literature, there is some debate as to whether the disposition effect could be caused by discrepant biases, such as motivated beliefs (Heinke et al.,, 2023), the gambler’s fallacy (Jiao,, 2017) or general underinference (Pitkäjärvi,, 2022). More examples range from linking confirmation bias to several stylized facts in financial markets (Pouget et al.,, 2017); or confidence biases to poor investment performance (Ahmad and Shah,, 2020) and biased memory (Huffman et al.,, 2022). As previously mentioned, utilizing an approach that is able to tell conflicting biases apart has the advantage to better pinpoint what biases are precisely underpinning certain behavioral regularities; or even let us know if some biases are simply an artifact of using a less complete model.
Third, the problem of conflicting biases also arises in the field of behavioral interventions. Boosting is a psychological technique aimed at de-biasing individuals when they suffer some kind of cognitive bias. One of the challenges mentioned in this area of research is the fact that these interventions normally target one bias at a time, when indeed multiple biases could be at stake (Kahneman et al.,, 2021). Once again, having a method to distinguish between different conflicting biases could come in handy in order to focus the intervention on those biases which end up being more common, or drive inference the most.
@@
The paper continues as follows: Section 2 explains the theoretical framework, the updating setting and progressively introduces different biases into the model. Section 3 deals with the experimental design and presents the belief-measuring tool. In section 4, I bring the experimental data and the theory together and compare two regression models, which differ in the amount of biases they incorporate. Section 5 discusses the results of the experiment, and section 6 concludes with some final remarks.
@@ 2.1 Theoretical framework
Let an agent observe a signal $S=\{s_{1},...,s_{n}\}$ consisting of the realization of $n$ independent and identically distributed Bernoulli trials. Let $p$ denote the probability of success of each trial ($s_{i}=1$) and $1-p$ denote the probability of each failure ($s_{i}=0$). For such data generating process (DGP), the likelihood function is the probability mass function of a binomial distribution with parameters $(n,p)$:
@@
where $k={\sum_{i=1}^{n}s_{i}}$ and $(n-k)$ are the number of successes and failures in the DGP respectively.
Let $\Omega=(0,1)$ be the set of possible values that $p$ may take, [^4] The state-space set $\Omega$ can have two different interpretations. One could either consider an agent who is not uncertain about her own beliefs, or one who is indeed uncertain about her own beliefs (à la cognitive uncertainty (Enke and Graeber,, 2023)). In the former case, the set $\Omega$ would specify the state-space of an agent who forms beliefs over every possible realisation of the objective parameter $p\in(0,1)$. In the latter case, the agent would form beliefs over a binary state-space (whether the signal $s_{i}$ takes value $1$ or 00), and the set $\Omega$ would, in this case, represent the subjective state-space ($p$), which describes the set of possible states where the agent expresses uncertainty about her own beliefs. Both alternatives are contemplated in this model. and let a prior belief $\pi(p)$ be beta distributed with prior parameters $(a_{0},b_{0})$. Namely,
@@
where $a_{0}$, $b_{0}$ $>0$, and $B(.)$ is the beta function.
Given a prior $\pi(p)$ and a likelihood $L(p|s_{1}...s_{n})$ the agent forms a posterior. As the beta distribution is a conjugate prior of the binomial distribution, a Bayesian agent updates her beliefs such that her posterior distribution of $p$, $\pi(p|a_{n},b_{n})$, is also beta distributed with parameters $a_{n},b_{n}$. This means that: [^5] See section A in the appendix for a proof; as well as a proof to derive eq.(4) and (5), which suffice to derive all other equations of posterior parameters in section 2.
@@
In order to incorporate updating biases, let a non-Bayesian agent use a distorted likelihood and prior when she updates. [^6] These distortions are often seen (Benjamin,, 2019) as part of an “as-if model.” This means that the model does not take the stand that biased agents actually follow Bayes’ Theorem with different likelihood and prior functions, but instead, that these distortions imply equivalent behavior to agents interpreting too little, or too much, from information signals (or prior beliefs) when they update. These distortions can be expressed as exponential deviations of the likelihood and prior. Namely, $\tilde{L}(p|s_{1},\dots,s_{n})=(L(p|s_{1},\dots,s_{n}))^{\gamma}$ would represent the distorted likelihood and $\tilde{\pi}(p)=(\pi(p))^{\delta}$ would represent a distorted prior of the non-Bayesian agent ($\gamma,\delta>0$). The parameters $\gamma\neq 1$ and $\delta\neq 1$ indicate deviations from Bayesian updating due to distortions of the likelihood and prior respectively. With these modified functions such a non-Bayesian agent has a posterior beta distribution with parameters $\tilde{a}_{n},\tilde{b}_{n}$ such that:
@@
In particular $\gamma>1$ indicates “overinference” while $\gamma<1$ shows “underinference”; that is, believing that the information signal $S$ is more/less informative than it actually is. Similarly, parameter $\delta>1$ indicates “base rate overuse” and $\delta<1$, “base rate neglect”; which implies that the prior is more/less informative than a Bayesian agent perceives it to be. This bias structure conceptually resembles Grether-regressions, [^7] That is, binary-state regressions where inference biases and base-rate biases are identified. as the same type of biases can be identified. The next section shows how equations (4) and (5) can be extended to incorporate more biases.

### v4

@@ Abstract
This paper proposes a unified theoretical model to identify and test a comprehensive set of probabilistic updating biases within a single framework. The model achieves separate identification by focusing on the updating of belief distributions, rather than point beliefs alone. Estimating the model in a laboratory experiment reveals significant individual heterogeneity: all tested biases are present and exhibit systematic co-occurrence patterns across individuals, with motivated-belief biases (optimism and pessimism) and sequence-related biases (gambler’s and hot-hand fallacy) emerging as key drivers of biased inference. At the population level most biases average out, but base-rate neglect remains a persistent influence. This study contributes to the belief-updating literature by providing a methodological toolkit for researchers examining links between conflicting biases and connections between updating biases and other behavioral phenomena.
00footnotetext: For their invaluable support, I would like to thank Elias Tsakas, Thomas Meissner and Matthias Wibral. For helpful comments, I thank three anonymous referees, Mohammed Abdellaoui, Miguel Ballester, En Hua Hu, Peiran Jiao, Emmanuel Kemel, Diego Marino-Fages, Antonio Penta, Arno Riedl, David Rojo Arjona, Michael Thaler, Severine Toussaert, and seminar participants at IMEBESS in Riga, ASFEE in Grenoble, Behavioural Game Theory Workshop in Norwich, Cognitive Foundations of Decision Making in Rabat, PhD conference in Micro Theory and Behavioral Economics in Berlin, and Maastricht University. I acknowledge funding from the Maastricht Graduate School of Business and Economics (GSBE) for primary data collection No. G.23.5168. Homepage: pgonzalezfernandez.com. E-mail: pedro.gonzalez-fernandez@awi.uni-heidelberg.de.
@@ 1 Introduction
Over the last decades, substantial effort has gone into identifying and categorizing belief-updating biases, both within Economics and Psychology (Edwards, 1968; Tversky and Kahneman, 1974; Grether, 1980). Scholars have typically studied these biases separately. As a result, it has become difficult to tell whether a person’s updating behavior should be attributed to one bias or another. Despite recent progress toward more comprehensive representations of biased updating (Benjamin, 2019; Stango and Zinman, 2023; Bordalo et al., 2023, 2025), there is still a lack of a unifying framework that can separate multiple potentially conflicting biases within the same empirical setting.
The basic problem is that different belief-updating biases may be observationally equivalent. Consider a person who appears to update “too much” after receiving new information. A traditional model may ask whether this reflects base-rate neglect, namely underweighting prior beliefs, or overinference, meaning placing too much weight on the new signal. But other biases may generate a similar updating pattern. If the person has preferences over outcomes, is she overreacting because she generally overinterprets information, or because she is optimistic or pessimistic about the state she prefers? Is she instead reacting strongly because the signal confirms her prior beliefs? Or does the same posterior reflect unusually concentrated beliefs, that is, overprecision? Once some biases are left outside the model, behavior may be attributed to the wrong one. In that sense, some biases may appear empirically important partly because other competing distortions are not being modeled.
@@
Distinguishing such confounded biases matters for at least two reasons. First, different biases can lead to very different actions even when they produce similar posteriors. For example, a person exhibiting confirmation bias may keep voting for the same political party after receiving social-media information that aligns with her previous beliefs, while someone who simply neglects prior information might be more likely to change her mind if reminded of the actual base rate. Second, once these biases are separately identified, one can ask which distortions are more prevalent, which matter most for inference, and which ones tend to appear together within individuals.
A central challenge in addressing this problem is methodological. The way beliefs are usually measured may not be rich enough to identify a broad set of potentially confounded biases. Much of the literature relies on point beliefs to study biased updating (Benjamin, 2019). While convenient, point beliefs offer limited flexibility to distinguish between different updating patterns. This paper proposes an alternative framework based on belief distributions.
@@
The paper constructs a structural model that allows multiple conflicting belief-updating biases to be separately identified using prior and posterior belief distributions. [^1] While Jiao et al. (2020) use belief distributions to incorporate several biases simultaneously, their framework is not well suited to probabilistic biases, relies on normally distributed beliefs, incorporates a more limited set of distortions, and is not taken to experimental data. Belief distributions can be understood either as beliefs over multiple outcomes or as uncertainty over one’s own beliefs when the underlying state space is binary, in a sense related to cognitive uncertainty (Enke and Graeber, 2023). In either interpretation, the key idea is that belief distributions provide a sharper language for defining and distinguishing belief biases.
@@
The theoretical framework considers an agent who observes a sequence of Bernoulli trials. Prior beliefs are assumed to follow a beta distribution and, under Bayesian updating, posteriors remain beta distributed because the beta distribution is conjugate to the binomial distribution. [^2] This updating environment is comparable to settings used in the ambiguity-aversion literature (see Abdellaoui et al. (2025) for an example) and in work on cognitive uncertainty (Enke and Graeber, 2023). To introduce biases, the model allows for distortions of the likelihood and the prior. These distortions generate non-Bayesian posterior beta distributions and make it possible to identify core biases such as over- and underinference (Khaw et al., 2021; Ba et al., 2022; Augenblick et al., 2025b) and base-rate neglect or overuse (Benjamin et al., 2019; Bucher and Glimcher, 2022; Enke et al., 2023). The framework also accommodates asymmetries in reactions to good and bad news, preference-based biases (Eil and Rao, 2011; Zimmermann, 2020; Möbius et al., 2022), confirmation bias (Rabin and Schrag, 1999; Charness and Dave, 2017; Zhenxun, 2024), sequence-related biases such as the hot-hand and gambler’s fallacies (Rabin, 2002; Rabin and Vayanos, 2010; Prat-Carrabin et al., 2024), and precision distortions (Moore et al., 2015; Augenblick et al., 2025a) that affect the variance of posterior beliefs.
@@
The model is estimated using data from a novel laboratory experiment. Participants solve a series of urn tasks in which they must guess the percentage of red balls in a selected urn drawn from a pool of 99 urns with varying red-blue compositions. They receive information through sequences of ball draws from the selected urn and report their belief beta distributions twice: once after an initial signal, and again after observing a second sequence of independently drawn signals from the same urn.
The results support the idea that richer models of biased updating reduce the apparent importance of some biases found in more parsimonious specifications. At the population level, overinference and base-rate neglect are both significant when the model allows only these distortions. Once the model incorporates a broader set of biases, however, the evidence for overinference disappears and base-rate neglect remains as the only distortion at the aggregate level. This aggregate bias distribution is also broadly consistent with the classic reduced-form belief-updating literature when the same environment is translated into a traditional binary-state specification.
At the same time, aggregate results mask substantial individual heterogeneity. [^3] The importance of individual heterogeneity for biased updating has recently been emphasized by Khaw et al. (2021) and Alós-Ferrer and Garagnani (2023). Their work supports the idea that average reports may look almost Bayesian or noisy, while individual-level data reveal systematic but heterogeneous deviations that partly cancel out in the aggregate. The individual-level analysis is therefore more informative. First, all tested biases are present to some extent in the data once the full array of biases is taken into account, revealing a much richer pattern of distorted updating than is visible in the aggregate. This also makes it possible to identify distorted behavior that would otherwise be statistically indistinguishable from Bayesian updating or too noisy to classify. Second, motivated-belief biases (optimism and pessimism) and sequence-related biases (the hot-hand and gambler’s fallacies) emerge as the main drivers of biased inference, with the hot-hand fallacy standing out as the most commonly exhibited distortion, while confirmatory biases are comparatively scarce. Finally, biases do not appear independently. Instead, they exhibit systematic co-occurrence patterns across individuals. Once again, sequence-related biases are central: A group of distortions (especially overprecision, overinference, and motivated-beliefs in general) seem to cluster around the hot-hand fallacy, while a different class of biases (such as prior-based distortions) tend to be jointly exhibited with the gambler’s fallacy.
@@
The paper contributes to the literature in three ways. First, it contributes to recent work that seeks to understand the links between different updating distortions. For example, Heger and Papageorge (2018) and Gneezy et al. (2023) study how wishful thinking can affect overconfidence; Charness and Dave (2017) and Zhenxun (2024, 2023) distinguish behavior stemming from motivated beliefs and unmotivated confirmation bias; Aydogan et al. (2025) build on Rabin and Schrag (1999) to separate conservatism from confirmation bias; and Chopra et al. (2024) study the trade-off between accuracy concerns and belief-confirmation motives in political information processing. These are important steps toward understanding how biases interact. More generally, however, in the absence of a framework that accommodates a broad range of distortions simultaneously, empirical links between biases may themselves be confounded. Bias $A$ may appear related to bias $B$ only because a third bias $C$ is omitted.
@@
Second, the paper contributes to research that seeks belief-based explanations for behavioral phenomena in settings where several competing distortions are plausible. Political polarization, for instance, has been explained from the perspective of overconfidence (Ortoleva and Snowberg, 2015), which is closely related to overprecision, and also from the perspective of confirmation bias (Del Vicario et al., 2017). In finance, the disposition effect has been linked to motivated beliefs (Heinke et al., 2023), the gambler’s fallacy (Jiao, 2017), and general underinference (Pitkäjärvi, 2022). Other examples include work linking confirmation bias to stylized facts in financial markets (Pouget et al., 2017), or confidence distortions to poor investment performance (Ahmad and Shah, 2020) and biased memory (Huffman et al., 2022). A framework that distinguishes conflicting biases more cleanly can therefore help identify which distortions are actually driving particular behavioral regularities, and whether some candidate explanations are partly artifacts of less complete models.
Third, the paper contributes to the literature on behavioral interventions. Boosting refers to de-biasing techniques that aim to improve decision-making when individuals suffer from cognitive distortions. One challenge emphasized in this literature is that interventions often target one bias at a time, even though several may be jointly at work (Kahneman et al., 2021). A method that distinguishes between multiple conflicting belief-updating biases can therefore help direct interventions toward the distortions that are most prevalent or that matter most for inference.
@@
The remainder of the paper is organized as follows. Section 2 introduces the theoretical framework and progressively incorporates different biases into the model. Section 3 presents the experimental design and the belief-measurement tool. Section 4 brings the theoretical and experimental components together and compares two regression models that differ in the number of biases they incorporate. Section 5 discusses the results, and Section 6 concludes.
@@ 2.1 Theoretical framework
Let an agent observe a signal $S=\{s_{1},...,s_{n}\}$ consisting of the realization of $n$ independent and identically distributed Bernoulli trials. Let $p$ denote the probability of success of each trial ($s_{i}=1$) and 1p1-p denote the probability of each failure ($s_{i}=0$). For such data generating process (DGP), the likelihood function is the probability mass function of a binomial distribution with parameters $(n,p)$:
@@
where $k={\sum_{i=1}^{n}s_{i}}$ and (nk)(n-k) are the number of successes and failures in the DGP respectively.
Let $\Omega=(0,1)$ be the set of possible values that $p$ may take, [^4] The state-space set $\Omega$ can have two different interpretations. One could either consider an agent who is not uncertain about her own beliefs, or one who is indeed uncertain about her own beliefs (à la cognitive uncertainty (Enke and Graeber, 2023)). In the former case, the set $\Omega$ would specify the state-space of an agent who forms beliefs over every possible realization of the objective parameter $p\in(0,1)$. In the latter case, the agent would form beliefs over a binary state-space (whether the signal $s_{i}$ takes value $1$ or 0), and the set $\Omega$ would, in this case, represent the subjective state-space ($p$), which describes the set of possible states where the agent expresses uncertainty about her own beliefs. The model is agnostic to either one of these interpretations. and let a prior belief $\pi(p)$ be beta distributed with prior parameters $(a_{0},b_{0})$. Namely,
@@
where $a_{0}$, $b_{0}$ >0>0, and $B(.)$ is the beta function.
Given a prior $\pi(p)$ and a likelihood $L(p|s_{1}...s_{n})$ the agent forms a posterior. As the beta distribution is a conjugate prior of the binomial distribution, a Bayesian agent updates her beliefs such that her posterior distribution of $p$, $\pi(p|a_{n},b_{n})$, is also beta distributed with parameters $a_{n},b_{n}$. This means that: [^5] See Appendix A for a proof; as well as a proof to derive eq.(4) and (5), which suffice to derive all other equations of posterior parameters in section 2.
@@
In order to incorporate updating biases, let a non-Bayesian agent use a distorted likelihood and prior when she updates. [^6] These distortions are often seen (Benjamin, 2019) as part of an “as-if model.” This means that the model does not take the stand that biased agents actually follow Bayes’ Theorem with different likelihood and prior functions, but instead, that these distortions imply equivalent behavior to agents interpreting too little, or too much, from information signals (or prior beliefs) when they update. These distortions can be expressed as exponential deviations of the likelihood and prior. Namely, $\tilde{L}(p|s_{1},\dots,s_{n})=(L(p|s_{1},\dots,s_{n}))^{\gamma}$ would represent the distorted likelihood and $\tilde{\pi}(p)=(\pi(p))^{\delta}$ would represent a distorted prior of the non-Bayesian agent ($\gamma,\delta>0$). The parameters $\gamma\neq 1$ and $\delta\neq 1$ indicate deviations from Bayesian updating due to distortions of the likelihood and prior respectively. With these modified functions such a non-Bayesian agent has a posterior beta distribution with parameters $\tilde{a}_{n},\tilde{b}_{n}$ such that:
@@
In particular $\gamma>1$ indicates “overinference” while $\gamma\in[0,1)$ shows “underinference”; that is, believing that the information signal $S$ is more/less informative than it actually is. A parameter of $\gamma<0$ reflects updating against the information signal. Similarly, parameter $\delta>1$ indicates “base-rate overuse” and $\delta<1$, “base-rate neglect”; which implies that the prior is more/less informative than a Bayesian agent perceives it to be. This bias structure conceptually resembles Grether-regressions, [^7] That is, binary-state regressions where inference biases and base-rate biases are identified. as the same type of biases can be identified. The next section shows how equations (4) and (5) can be extended to incorporate incrementally more biases.